# Turning Sand to Gold: Recycling Data to Bridge On-Policy and Off-Policy Learning via Causal Bound

### v1

@@ 5.2 Controlled Comparison with Baseline Agents
We evaluate the impact of incorporating the SUFT OPE term by comparing agents that include our proposed term to baseline agents without it, ensuring both types share identical architecture and hyperparameters where the presence of our proposed term is the only difference between them. This controlled experiment is designed to isolate the influence of our method, ensuring fair comparisons.
In addition, we conduct a $t$-test for each environment against a two-sided alternative using 10 different random seed runs, where the null hypothesis is the baseline result. The $p$-values provide robust statistical significance results for evaluating the effectiveness of our proposed method. The complete experimental results, a detailed ablation study of our method against other components, and the evaluation formulas are detailed in Appendix C. We also analyze the computational and space overhead introduced by our method, showing it is negligible as detailed in Appendix C.
@@
For the Atari benchmark, SUFT demonstrates significant improvements for all of the agents. SUFT surpasses the baseline agent in 52, 48, and 50 games out of the 57 games, with statistically significant gains ($p$-value $<$ 0.05) in 49, 37, and 40 games. These results are for the DQN, Double DQN, and PPO agents, respectively.
We also analyze the results for the valid Atari games (according to the evaluation formula in Appendix C). For the DQN agent, SUFT outperforms the baseline agent in 30 out of the 32 valid environments, achieving a reward improvement of over 100% in 50% of them and over 10% in 90.6% of them, as shown in Appendix C. For the Double DQN agent, SUFT outperforms the baseline agent in 36 out of the 42 valid environments, achieving a reward improvement of over 100% in 21.4% of them and over 10% in 71.4% of them, as shown in Figure 1 (Left). For the PPO agent, SUFT surpasses the baseline agent in 40 out of the 43 valid environments, achieving a reward improvement of over 100% in 39.5% of them and over 10% in 83.7% of them, as shown in Figure 1 (Right).
Additionally, when evaluating the mean reward ratio (see Appendix C for the formula details), SUFT demonstrates profound gains relative to the baseline agents. For the DQN, Vanilla DQN, Double DQN, and PPO agents, SUFT achieves a mean reward ratio of 2,427.06%, 230.49%, 260.03%, and 203.21%, respectively, as shown in Figure 2.
Moreover, in several Atari games, SUFT not only surpasses the baseline agent but also achieves higher rewards than an average human, as shown in Figure 3. This achievement highlights the potential of our method to boost agents’ performance even above the human level, while keeping constrained environment interactions and resource demands. The human rewards are sourced from [79].
@@
For the MuJoCo benchmark, SUFT demonstrates improvements in the SAC and PPO agents. For the SAC agent, SUFT surpasses the baseline agent in 4 out of 5 environments, with statistically significant gains in 3 environments. For the PPO agent, SUFT surpasses the baseline agent in 4 out of 5 environments (in the other environment, the score is the same), with statistically significant gains in 2 environments. In regard to the mean reward ratio, SUFT achieves 224.53% for the SAC agent and 111.21% for the PPO agent, as shown in Figure 2.
@@ 5.3 Buffer Size Reduction
We also evaluate the sample efficiency of SUFT by comparing its performance to an identical DQN agent utilizing a larger 100K experience replay buffer size. Remarkably, the SUFT agent achieves a mean reward ratio gain of 503.36%, outperforming an agent with a 25 times larger buffer size.
These results underscore the potential of our approach to obtain high-performance DRL agents with significantly smaller buffers, paving the way for more computationally efficient and accessible DRL methods by enhancing sample efficiency at negligible cost.
@@ 6 Conclusion
This paper presents SUFT, a causal upper-bound loss optimization method that enhances DRL sample efficiency and reduces computational demands at negligible cost. We begin by establishing a provable bound on the factual loss within the Neyman-Rubin potential outcomes framework and seamlessly adapting it to the DRL framework. Showing that on-policy loss is bounded by every agent’s standard off-policy loss and an additional SUFT OPE term. This term is computable by storing past value network outputs in the experience replay buffer, thereby reusing discarded data to improve agents’ performance. Experimental results across diverse environments and agents show that our method leads to significant improvements in convergence rates and reward outcomes.
@@ B.1 Definitions and Proofs
Inequality 1 is by $x\leq|x|$, equality 2 is by the definition of ${L_{1}}$ loss, inequality 3 is by reverse triangle inequality $||x_{1}|-|x_{2}||\leq|x_{1}-x_{2}|$, inequality 4 is by triangle inequality $|x_{1}+x_{2}|\leq|x_{1}|+|x_{2}|$, equality 5 is by absolute symmetry $|a|=|-a|$, and the definition of ${L_{1}}$ loss. ∎
@@
We assume a joint distribution $P(X,T,Y_{0},Y_{1})$ with the following components: $q_{t}:=\Pr(T=t)$ is the treatment probability. $P^{t}_{X}:=P(X|T=t)$ is the conditional distribution of observations. $P^{x}_{Y_{1}}:=P(Y_{1}|X=x)$ is the conditional distribution of potential outcomes given the target treatment, and $P^{x}_{Y_{0}}:=P(Y_{0}|X=x)$ is the conditional distribution of potential outcomes given the control treatment. We also assume that there are no hidden confounders in the observational data, which makes the estimated treatment effect identifiable. This is formalized by assuming that the standard strong ignorability condition holds: $(Y_{0},Y_{1})\mathrel{\perp\!\!\!\perp}T|X$, and $10<p(t=1|x)<$ 1 for all $x$. Strong ignorability is a sufficient condition for the estimated treatment effect to be identifiable [70, 57].
@@ B.2 Causal Bound Theorem
[…]
@@
$=q_{1}\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\ell_{\phi}(x,1)]+q_{0}\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\ell_{\phi}(x,0)]-q_{0}\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\ell_{\phi}(x,1)]-q_{1}\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\ell_{\phi}(x,0)]$ (7)
@@
Now we will use the linearity of expectation and Lemma B.11:
@@
Equality 6 is by Definitions B.6 and B.8, equality 7 is by Definitions B.5 and B.7, equality 8 is by Definition B.4, equality 9 is by linearity of expectations, equality 10 is by linearity of expectations and by Lemma B.11, inequality 11 is by B.1, equality 12 is by linearity of expectations and because $\phi$ is a constant on the expectation of $Y_{0},Y_{1}$, equality 13 is by Definition B.9 and $\delta^{1},\delta^{0}$ notation, equality 14 is by Definition B.10 and $\delta$ notation. ∎
@@ B.3 Causal Bound for $N$ Control Treatments
We now adjust the definitions and Theorem B.12 proof from the binary treatment scenario to a single target treatment and N control treatments scenario.
[…]
We assume joint distribution $P(X,T,Y_{1},Y_{2},...,Y_{N+1})$ with the following components: $q_{t}:=\Pr(T=t)$ is the treatment probability. $P^{t}_{X}:=P(X|T=t)$ is the conditional distribution of observations. $P^{x}_{Y_{1}}:=P(Y_{1}|X=x)$ is the conditional distribution of potential outcomes given the target treatment, and $\forall j\in\{2,...,N+1\}$ $P^{x}_{Y_{j}}:=P(Y_{j}|X=x)$ is the conditional distribution of potential outcomes given a control treatment $t=j$.
@@
We now generalize Theorem B.12 from a single control treatment scenario to the $N$ control treatments scenario as follows:
@@
As we prove in Theorem B.12 where the control treatment is $t=2$ instead of $t=0$.
@@
$\sum_{j=2}^{N}\left(\frac{1}{N-1}q_{1}\epsilon^{1}_{F_{\phi}}+q_{j}\epsilon^{j}_{F_{\phi}}-q_{j}\epsilon^{1,j}_{CF_{\phi}}-\frac{1}{N-1}q_{1}\epsilon^{j}_{CF_{\phi}}\right)\leq\sum_{j=2}^{N}\left(q_{j}\psi^{1,j}_{\phi}+\frac{1}{N-1}q_{1}\psi^{j}_{\phi}+q_{j}\delta^{1,j}+\frac{1}{N-1}q_{1}\delta^{j}\right)$
@@
$\sum_{j=2}^{N+1}\left(\frac{1}{N}q_{1}\epsilon^{1}_{F_{\phi}}+q_{j}\epsilon^{j}_{F_{\phi}}-q_{j}\epsilon^{1,j}_{CF_{\phi}}-\frac{1}{N}q_{1}\epsilon^{j}_{CF_{\phi}}\right)$ $=\sum_{j=2}^{N}\left(\frac{1}{N}q_{1}\epsilon^{1}_{F_{\phi}}+q_{j}\epsilon^{j}_{F_{\phi}}-q_{j}\epsilon^{1,j}_{CF_{\phi}}-\frac{1}{N}q_{1}\epsilon^{j}_{CF_{\phi}}\right)$
$+\frac{1}{N}q_{1}\epsilon^{1}_{F_{\phi}}+q_{N+1}\epsilon^{N+1}_{F_{\phi}}-q_{N+1}\epsilon^{1,N+1}_{CF_{\phi}}-\frac{1}{N}q_{1}\epsilon^{N+1}_{CF_{\phi}}$
@@
$+\frac{1}{N}q_{1}\epsilon^{1}_{F_{\phi}}+q_{N+1}\epsilon^{N+1}_{F_{\phi}}-q_{N+1}\epsilon^{1,N+1}_{CF_{\phi}}-\frac{1}{N}q_{1}\epsilon^{N+1}_{CF_{\phi}}$
@@
$+q_{N+1}\psi^{1,N+1}_{\phi}+\frac{1}{N}q_{1}\psi^{N+1}_{\phi}+q_{N+1}\delta^{1,N+1}+\frac{1}{N}q_{1}\delta^{N+1}$
@@
Inequality 15 is by the inductive hypothesis, Inequality 16 is from the proof of Theorem B.12 where the control treatment is $t=N+1$ instead of $t=0$. ∎
@@ B.4 Reduction from Causal to DRL Framework
Definition B.4 in the DRL framework refers to the temporal difference loss between the accumulated rewards and the Q-value estimation, and it is defined as follows:
@@
Definitions B.5 and B.13 in the DRL framework assigns the temporal difference loss with experiences collected on-policy, with identical target and behavior policies from $D_{1}$, and it is defined as follows:
@@
Definitions B.7 and B.15 in the DRL framework assigns the temporal difference loss with experiences gathered off-policy, with different target and behavior policies from $D_{0}$, and it is defined as follows:
@@
Definitions B.9 and B.17 interpretation in the DRL framework is an OPE term, and it is defined as follows:
@@
Theorems B.12 and B.19 interpretation in the DRL framework establishes an upper bound to the on-policy loss, combining the off-policy loss and the SUFT OPE term.
@@
Definition B.4 in the DRL framework refers to the temporal difference loss between the accumulated rewards and the V estimation, and it is defined as follows:
@@
Definitions B.5 and B.13 in the DRL framework assigns the temporal difference loss with experiences collected on-policy, with identical target and behavior policies from $D_{1}$, and it is defined as follows:
@@
Definitions B.7 and B.15 in the DRL framework assigns the temporal difference loss with experiences gathered off-policy, with different target and behavior policies from $D_{0}$, and it is defined as follows:
@@
Definitions B.9 and B.17 interpretation in the DRL framework is an OPE term, and it is defined as follows:
@@
Theorems B.12 and B.19 interpretation in the DRL framework establishes an upper bound to the on-policy loss, combining the off-policy loss and the SUFT OPE term.
@@
Note that since the constant term, $\delta$, is independent of $V$, it is disregarded during the gradient-based optimization of $V$, so the SUFT causal bound objective term is as follows:
@@
An important alignment needs to be made to address a difference between the static observational data assumed by the causal framework and the dynamic experience replay buffer in the online DRL framework. In the causal framework, all observations are gathered once, so the data is fixed and irrevocably labeled either to be from the target treatment or from the control treatment in the binary treatment case. On the other hand, in the online DRL framework, the observational data changes over time, as we continuously populate the experience replay buffer and update the target policy. On each policy optimization step, all of the samples stored in the buffer that were labeled as target policy are turned into samples that are labeled as a control policy. This phenomenon of changing the observational data labels when updating the policy in the DRL framework is a gap between the causal and DRL frameworks. In the on-policy case, the experience replay buffer contains only on-policy observations, meaning that $q_{1}=1$. When updating the target policy, this data becomes off-policy, and the data that was previously labeled as target policy turns into control policy, shifting the samples’ labels together with the treatment probability $q_{0_{\text{new}}}=q_{1_{\text{old}}}=1$.

### v2

@@ 5.2 Controlled Comparison with Baseline Agents
We evaluate the impact of incorporating the SUFT OPE term by comparing agents that include our proposed term to baseline agents without it, ensuring both types share identical architecture and hyperparameters, where the presence of our proposed term is the only difference between them. This controlled experiment is designed to isolate the influence of our method, ensuring fair comparisons.
In addition, we conduct a $t$-test for each environment against a two-sided alternative using 10 different random seed runs, where the null hypothesis is the baseline result. The $p$-values provide robust statistical significance results for evaluating the effectiveness of our proposed method.
The complete experimental results, the evaluation formulas, and more empirical analysis of our method are detailed in Appendix˜D.
@@
On the Atari benchmark, SUFT demonstrates significant improvements for all of the agents. On the DQN, Double DQN, and PPO agents, SUFT surpasses the baseline agent in 52, 48, and 50 games out of the 57 games, respectively, with statistically significant gains ($p$-value << 0.05) in 49, 37, and 40 games.
We also analyze the results for the valid Atari games (according to the evaluation formula in Appendix˜D). For the DQN agent, SUFT outperforms the baseline agent in 28 out of the 29 valid games, achieving a reward improvement of over 100% in 48.3% of them and over 10% in 93.1% of them, as shown in Figure˜11. For the Double DQN agent, SUFT outperforms the baseline agent in 35 out of the 40 valid games, achieving a reward improvement of over 100% in 20% of them and over 10% in 72.5% of them, as shown in Figure˜1 (Left). For the PPO agent, SUFT surpasses the baseline agent in 39 out of the 42 valid games, achieving a reward improvement of over 100% in 38.1% of them and over 10% in 83.3% of them, as shown in Figure˜1 (Right).
Additionally, when evaluating the mean reward ratio, SUFT demonstrates profound gains relative to the baseline agents. For the DQN, Vanilla DQN, Double DQN, and PPO agents, SUFT achieves a mean reward ratio of 383.19%, 228.12%, 226.91%, and 197.34%, respectively, as shown in Figure˜2.
Moreover, in several Atari games, SUFT not only surpasses the baseline agent but also achieves higher rewards than an average human, as shown in Figure˜3. This achievement highlights the potential of our method to boost agents’ performance even above the human level, while keeping constrained environment interactions and resource demands. The human rewards are sourced from [79].
@@
For the MuJoCo benchmark, SUFT demonstrates improvements in the SAC and PPO agents. For the SAC agent, SUFT surpasses the baseline agent in 4 out of 5 environments, with statistically significant gains in 3 environments. For the PPO agent, SUFT surpasses the baseline agent in 4 out of 5 environments (in the other environment, the score is the same), with statistically significant gains in 2 environments. With regard to the mean reward ratio, SUFT achieves 224.53% for the SAC agent and 111.21% for the PPO agent, as shown in Figure˜2.
@@ 5.3 Buffer Size Reduction
We also evaluate the sample efficiency of SUFT by comparing its performance to an identical DQN agent utilizing a larger 100K experience replay buffer size. Remarkably, SUFT achieves a mean reward ratio gain of 437.05%, outperforming an identical agent that uses a 25 times larger buffer size.
These results underscore the potential of our approach to obtain high-performance DRL agents with significantly smaller buffers, paving the way for more computationally efficient and accessible DRL methods by enhancing sample efficiency at a negligible cost.
5.4 Ablation Study
We conduct an ablation study to isolate the contribution of the SUFT OPE term to DQN performance. As shown Table˜1, we compare SUFT DQN against several alternatives: increasing the buffer size, a Vanilla DQN, and a Double DQN. While enlarging the buffer or adopting more complex DQN variants leads to improvements, incorporating the SUFT OPE term yields the highest Human Normalized Mean of 63.45%, outperforming all other baselines that require higher resources and computational costs. These results highlight the potential of our method to improve the agent’s performance by enhancing sample efficiency at a negligible cost, providing a simple yet highly effective method for improving agent performance at a negligible cost.
Table 1: An ablation study on the DQN agent comparing the SUFT additional term impact against enlarging the buffer size, adding a Vanilla DQN, and adding a Double DQN. The DQN, Vanilla DQN, Double DQN, and SUFT DQN use a 4K buffer size. All agents are using $L_{2}$} loss.
Agent Human Normalized Mean (%)
DQN 25.04%
DQN, buffer 100K 29.29%
Vanilla DQN 41.26%
Double DQN 42.31%
SUFT DQN 63.45%
5.5 Causal Bound Optimization vs Inertia Regularizer
To isolate the effect of our causal mechanism and ensure that the significant improvements of our method stem from the causal bound optimization rather than merely from a regularization effect, we conduct a controlled experiment. Specifically, we evaluate a double DQN agent that adds an inertia regularizer to the standard temporal difference loss function in the form of:
$\left(Q_{\text{current network}}-Q_{\text{target network}}\right)^{2}$
Where $Q_{\text{target network}}$ denotes the slowly updated target network, as done in a standard double DQN. The experiment uses the same configurations as both the SUFT double DQN and the baseline double DQN, with the additional regularization term being the only difference from the baseline agent. The results in Table˜11 show that the inertia regularization significantly degrades double DQN performance. Resulting in an 83.21% mean reward ratio. This provides evidence that the gains from our method, which result in a 226.91% mean reward ratio improvement, arise from the causal bound optimization, where the old Q-network values used for the OPE term correspond to the behavior policy that generated the experiences, thereby mitigating the divergence between the on-policy and off-policy methods. These results indicate that our method functions not merely as a stability-promoting regularizer, but as a principled causal bound optimization.
@@ 6 Conclusion
This paper presents SUFT, a causal upper-bound loss optimization method that enhances DRL sample efficiency and reduces computational demands at a negligible cost. We begin by establishing a provable bound on the factual loss within the Neyman-Rubin potential outcomes framework and seamlessly adapting it to the DRL framework, showing that on-policy loss is bounded by every agent’s standard off-policy loss and an additional SUFT OPE term. This term is computable by storing past value network outputs in the experience replay buffer, thereby reusing discarded data to improve agents’ performance. Experimental results across diverse environments and agents show that our method leads to significant improvements in convergence rates and reward outcomes. Looking ahead to future work, we see physical robotics as a particularly promising domain, where SUFT’s reduced-resource method could make DRL more practical under real-world constraints.
@@ B.1 Definitions and Proofs
Inequality 1 is by $x\leq|x|$, equality 2 is by the definition of $L_{1}$} loss, inequality 3 is by reverse triangle inequality $||x_{1}|-|x_{2}||\leq|x_{1}-x_{2}|$, inequality 4 is by triangle inequality $|x_{1}+x_{2}|\leq|x_{1}|+|x_{2}|$, equality 5 is by absolute symmetry |a|=|a||a|=|-a|, and the definition of $L_{1}$} loss. ∎
@@
We assume a joint distribution $P(X,T,Y_{0},Y_{1})$ with the following components: $q_{t}:=\Pr(T=t)$ is the treatment probability. $P^{t}_{X}:=P(X|T=t)$ is the conditional distribution of observations. $P^{x}_{Y_{1}}:=P(Y_{1}|X=x)$ is the conditional distribution of potential outcomes given the target treatment, and $P^{x}_{Y_{0}}:=P(Y_{0}|X=x)$ is the conditional distribution of potential outcomes given the control treatment. We also assume that there are no hidden confounders in the observational data, which makes the estimated treatment effect identifiable. This is formalized by assuming that the standard strong ignorability condition holds: $(Y_{0},Y_{1})\mathrel{\perp\!\!\!\perp}T|X$, and $0<p(t=1|x)<1$ for all $x$. Strong ignorability is a sufficient condition for the estimated treatment effect to be identifiable [70, 57].
@@ B.2 Causal Bound Theorem
$\mathbb{E}_{y_{1}\sim P^{x}_{Y_{1}}}[\mathbb{E}_{y_{0}\sim P^{x}_{Y_{0}}}[\mathbf{L}(y_{1},\phi(x;\theta_{1}))]]$ $=\int_{\mathcal{Y}}\left(\int_{\mathcal{Y}}\mathbf{L}(y_{1},\phi(x;\theta_{1}))\cdot p(y_{0}|x)dy_{0}\right)\cdot p(y_{1}|x)dy_{1}$
$=\int_{\mathcal{Y}}\mathbf{L}(y_{1},\phi(x;\theta_{1}))\cdot\left(\int_{\mathcal{Y}}p(y_{0}|x)dy_{0}\right)\cdot p(y_{1}|x)dy_{1}$
$=\int_{\mathcal{Y}}\mathbf{L}(y_{1},\phi(x;\theta_{1}))\cdot p(y_{1}|x)dy_{1}$
@@
$=q_{1}\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\ell_{\phi}(x,1)]+q_{0}\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\ell_{\phi}(x,0)]$ (7)
$-q_{0}\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\ell_{\phi}(x,1)]-q_{1}\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\ell_{\phi}(x,0)]$
@@
Now we will use the linearity of expectation and Lemma˜B.11:
@@
Equality 6 is by Definitions˜B.6 and B.8, equality 7 is by Definitions˜B.5 and B.7, equality 8 is by Definition˜B.4, equality 9 is by linearity of expectations, equality 10 is by linearity of expectations and by Lemma˜B.11, inequality 11 is by ˜B.1, equality 12 is by linearity of expectations and because $\phi$ is a constant on the expectation of $Y_{0},Y_{1}$, equality 13 is by Definition˜B.9 and $\delta^{1},\delta^{0}$ notation, equality 14 is by Definition˜B.10 and $\delta$ notation. ∎
@@ B.3 Causal Bound for $N$ Control Treatments
We now adjust the definitions and Theorem˜B.12 proof from the binary treatment scenario to a single target treatment and N control treatments scenario.
We first set the definitions of the Neyman-Rubin potential outcome framework for the $N$ control treatments scenario: $\mathcal{T}=\{1,\ldots,N+1\}$ is the treatments space, where $t=j:\ 2\leq j\leq N+1$ is the control treatments, and $t=1$ is the target treatment compared against each control treatment. Under this framework, there are two types of potential outcomes depending on the treatment assignment: if $t=1$, the observed outcome $y=Y_{1}$ is from the target treatment, and $\forall j\in\{2,\ldots,N+1\}$ if $t=j$, the observed outcome $y=Y_{j}$ is from a control treatment.
We assume joint distribution $P(X,T,Y_{1},Y_{2},\ldots,Y_{N+1})$ with the following components: $q_{t}:=\Pr(T=t)$ is the treatment probability. $P^{t}_{X}:=P(X|T=t)$ is the conditional distribution of observations. $P^{x}_{Y_{1}}:=P(Y_{1}|X=x)$ is the conditional distribution of potential outcomes given the target treatment, and $\forall j\in\{2,\ldots,N+1\}$ $P^{x}_{Y_{j}}:=P(Y_{j}|X=x)$ is the conditional distribution of potential outcomes given a control treatment $t=j$.
@@
We now generalize Theorem˜B.12 from a single control treatment scenario to the $N$ control treatments scenario as follows:
@@
As we prove in Theorem˜B.12 where the control treatment is $t=2$ instead of $t=0$.
@@
$\sum_{j=2}^{N}\left(\frac{1}{N-1}q_{1}\epsilon^{1}_{F_{\phi}}+q_{j}\epsilon^{j}_{F_{\phi}}-q_{j}\epsilon^{1,j}_{CF_{\phi}}-\frac{1}{N-1}q_{1}\epsilon^{j}_{CF_{\phi}}\right)$ $\leq$
$\sum_{j=2}^{N}\left(q_{j}\psi^{1,j}_{\phi}+\frac{1}{N-1}q_{1}\psi^{j}_{\phi}+q_{j}\delta^{1,j}+\frac{1}{N-1}q_{1}\delta^{j}\right)$
@@
$\sum_{j=2}^{N+1}\left(\frac{1}{N}q_{1}\epsilon^{1}_{F_{\phi}}+q_{j}\epsilon^{j}_{F_{\phi}}-q_{j}\epsilon^{1,j}_{CF_{\phi}}-\frac{1}{N}q_{1}\epsilon^{j}_{CF_{\phi}}\right)$ $=$
$\sum_{j=2}^{N}\left(\frac{1}{N}q_{1}\epsilon^{1}_{F_{\phi}}+q_{j}\epsilon^{j}_{F_{\phi}}-q_{j}\epsilon^{1,j}_{CF_{\phi}}-\frac{1}{N}q_{1}\epsilon^{j}_{CF_{\phi}}\right)$ $+\frac{1}{N}q_{1}\epsilon^{1}_{F_{\phi}}+q_{N+1}\epsilon^{N+1}_{F_{\phi}}$
$-q_{N+1}\epsilon^{1,N+1}_{CF_{\phi}}-\frac{1}{N}q_{1}\epsilon^{N+1}_{CF_{\phi}}$
@@
$+\frac{1}{N}q_{1}\epsilon^{1}_{F_{\phi}}+q_{N+1}\epsilon^{N+1}_{F_{\phi}}$
$-q_{N+1}\epsilon^{1,N+1}_{CF_{\phi}}-\frac{1}{N}q_{1}\epsilon^{N+1}_{CF_{\phi}}$
@@
$+q_{N+1}\psi^{1,N+1}_{\phi}+\frac{1}{N}q_{1}\psi^{N+1}_{\phi}$
$+q_{N+1}\delta^{1,N+1}+\frac{1}{N}q_{1}\delta^{N+1}$
@@
Inequality 15 is by the inductive hypothesis, Inequality 16 is from the proof of Theorem˜B.12 where the control treatment is $t=N+1$ instead of $t=0$. ∎
@@ B.4 Reduction from Causal to DRL Framework
Definition˜B.4 in the DRL framework refers to the temporal difference loss between the accumulated rewards and the Q-value estimation, and it is defined as follows:
@@
Definitions˜B.5 and B.13 in the DRL framework assigns the temporal difference loss with experiences collected on-policy, with identical target and behavior policies from $D_{1}$, and it is defined as follows:
@@
Definitions˜B.7 and B.15 in the DRL framework assigns the temporal difference loss with experiences gathered off-policy, with different target and behavior policies from $D_{0}$, and it is defined as follows:
@@
Definitions˜B.9 and B.17 interpretation in the DRL framework is an OPE term, and it is defined as follows:
@@
Theorems˜B.12 and B.19 interpretation in the DRL framework establishes an upper bound to the on-policy loss, combining the off-policy loss and the SUFT OPE term.
@@
Definition˜B.4 in the DRL framework refers to the temporal difference loss between the accumulated rewards and the V estimation, and it is defined as follows:
@@
Definitions˜B.5 and B.13 in the DRL framework assigns the temporal difference loss with experiences collected on-policy, with identical target and behavior policies from $D_{1}$, and it is defined as follows:
@@
Definitions˜B.7 and B.15 in the DRL framework assigns the temporal difference loss with experiences gathered off-policy, with different target and behavior policies from $D_{0}$, and it is defined as follows:
@@
Definitions˜B.9 and B.17 interpretation in the DRL framework is an OPE term, and it is defined as follows:
@@
Theorems˜B.12 and B.19 interpretation in the DRL framework establishes an upper bound to the on-policy loss, combining the off-policy loss and the SUFT OPE term.
@@
Note that since the constant term, $\delta$, is independent of $V$, it is disregarded during the gradient-based optimization of $V$. Thus, the SUFT causal bound objective term is as follows:
@@
An important alignment needs to be made to address a difference between the static observational data assumed by the causal framework and the dynamic experience replay buffer in the online DRL framework. In the causal framework, all observations are gathered once, so the data is fixed and irrevocably labeled either to be from the target treatment or from the control treatment in the binary treatment case. On the other hand, in the online DRL framework, the observational data changes over time, as we continuously populate the experience replay buffer and update the target policy. On each policy optimization step, all of the samples stored in the buffer that were labeled as the target policy are turned into samples that are labeled as a control policy. This phenomenon of changing the observational data labels when updating the policy in the DRL framework is a gap between the causal and DRL frameworks. In the on-policy case, the experience replay buffer contains only on-policy observations, meaning that $q_{1}=1$. When updating the target policy, this data becomes off-policy, and the data that was previously labeled as target policy turns into control policy, shifting the samples’ labels together with the treatment probability $q_{0_{\text{new}}}=q_{1_{\text{old}}}=1$.